# ExteriorExtensions: A package for Macaulay2

### v1

@@ Abstract.
We explain a Macaulay2 implementation of a construction, which appeared in [HolweckOeding22], of a graded algebra structure on the direct sum of a Lie algebra $\mathfrak{g}$ (typically $\mathfrak{sl}_{n}$) and a $\mathfrak{g}$-module (typically a subspace of an exterior algebra ${\textstyle\bigwedge^{\bullet}}\mathbb{C}^{n}$). We implement brackets, a Killing form, matrix representations of adjoint operators and ranks of blocks of their powers.
@@ 1. Introduction
The Lie algebra $\mathfrak{g}=\mathfrak{sl}_{n}(\mathbb{F})$ is a matrix algebra with bracket $[A,B]=AB-BA$ for traceless matrices $A,B\in\mathbb{F}^{n\times n}$. The package ExteriorExtensions extends this algebra to a larger one $\mathfrak{a}=\mathfrak{g}\oplus M$ for a $\mathfrak{g}$-module $M$, and at this time we have only implemented this for $M$ a sum of exterior powers of the underlying vector space $V=\mathbb{F}^{n}$. We say that the extension algebra has GJD (Jordan decomposition consistent with the $G$-action) if the $G=\text{Lie}(\mathfrak{g})$-action on $M$ is consistent with conjugation on adjoint operators $\operatorname{ad}_{T}\colon\mathfrak{a}\to\mathfrak{a}$ (see [HolweckOeding22]). The package provides functionality for the standard matrix representations of adjoint operators on this algebra, as well as a tabulation of the ranks of blocks of powers of adjoint operators. We also compute the Killing matrix, which may help in identifying the algebra.
@@ 1.1. Exterior algebras
The exterior algebra on a vector space $V$ is the direct sum of exterior powers ${\textstyle\bigwedge^{\bullet}}V=\bigoplus_{k\geq 0}{\textstyle\bigwedge^{k}}V$, with product given by concatenation, which respects a $\mathbb{Z}$-grading. The exterior algebra on $V$ is Artinian when $V$ is finite dimensional. However, we can obtain a $\mathbb{Z}_{n}$-grading if $n=\dim V$ by choosing a volume form $0\neq\Omega\in{\textstyle\bigwedge^{n}}V$ and replacing products $\varphi\wedge\psi$ with $\Omega(\Omega(\varphi)\wedge\Omega(\psi))$ if $\deg(\varphi)+\deg(\psi)>m$. Moreover, we note that one can obtain a (possibly) different grading other than the degree grading by declaring ${\textstyle\bigwedge^{j}}V$ to be the grade-1 piece, and defining the rest of the graded pieces inductively.
@@ 1.2. Extending the Lie algebra by appending an exterior algebra
Now we consider the graded vector space $\mathfrak{a}=\mathfrak{g}\oplus{\textstyle\bigwedge^{k}}V\oplus{\textstyle\bigwedge^{2k}}V\oplus\dots=\mathfrak{g}\oplus({\textstyle\bigwedge^{\bullet}}V\setminus{\textstyle\bigwedge^{0}}V)$, where we delete the scalars from the exterior algebra, and set $\mathfrak{a}_{0}=\mathfrak{g}$ and $\mathfrak{a}_{i}={\textstyle\bigwedge^{ik\mod n}}V$ for $1\leq i<n=\dim V$.
@@
Recall that a structure tensor $B\in\mathfrak{a}^{*}\otimes\mathfrak{a}^{*}\otimes\mathfrak{a}$ representing the bilinear product $[\;,\;]\colon\mathfrak{a}\times\mathfrak{a}\to\mathfrak{a}$ is equivalent information to the algebra structure on $\mathfrak{a}$. To respect the grading we must require that $B_{i,j,k}=0$ whenever $i+j\not\equiv k\mod n$. In [HolweckOeding22, Theorem 3.12] we showed that requiring GJD implies that there is an essentially unique structure tensor (product) on $\mathfrak{a}$. We implement that product (a method function in M2 called bracket) as follows.
1.2.1. The structure tensor / bracket
@@
For $B_{0,0,0}$ we have the standard skew-commuting bracket on $\mathfrak{sl}_{n}$, namely $[A,B]=AB-BA$. For $B_{0,j,j}$ (and for $B_{j,0,j}$) with $j>0$ we use the standard Lie algebra action of $\mathfrak{sl}_{n}$ on the fundamental representation ${\textstyle\bigwedge^{k_{j}}}\mathbb{C}^{n}$, namely for $A\in\mathfrak{sl}_{n}$ and $T\in{\textstyle\bigwedge^{k}}V$ we take $[A,T]:=A.T$, where . denotes the Lie algebra action. Similarly $[T,A]=-A.T$.
@@
$\begin{matrix}\mathfrak{a}_{i}\times\mathfrak{a}_{j}&\to&\mathfrak{a}_{0}\\ (S,T)&\mapsto&[S,T]=-k\pi\left(\sum_{p,q}\left(\frac{\partial S}{\partial e_{p}}\lrcorner\frac{\partial*T}{\partial e_{q}}\right)\right)\end{matrix}\quad\quad\text{when }i+j=n\text{ and }j>n/2.$
@@ 1.2.2. Adjoint Operators
The standard matrix representation of the adjoint operator on $\mathfrak{a}$ defined by $\operatorname{Ad}_{x}(y)=[x,y]$ is constructed by choosing an ordered basis of $\mathfrak{a}$, and expressing the results of $\operatorname{Ad}_{x}(y)$ for each basis vector $y$ in terms of the standard basis as the corresponding column. One caveat is that we have to choose a basis of $\mathfrak{sl}_{n}$, and there is not a "standard" choice for the basis of the Cartan subalgebra of traceless diagonal matrices. We choose a common one $\{E_{ii}-E_{i+1,i+1}\mid i=1\ldots n-1\}$. Then we have to solve a system of equations to find the coordinates for the cartan rather than just extracting the entries of the matrix.
Respecting the grading yields block structure given by $\operatorname{Ad}_{x}(y)=(B_{i,j,i+j\mod n}(y))_{i,j}$. The ranks of the blocks of $\operatorname{Ad}_{x}$ are invariants of $x$, as are the ranks of blocks of the powers of $\operatorname{Ad}_{x}$. We collect these ranks in a table and display them with a function prettyBlockRanks.
The Jacobi identity is stated via adjoint operators as $[\operatorname{Ad}_{x},\operatorname{Ad}_{y}]=\operatorname{Ad}_{[x,y]}$ [FultonHarris].

### v2

@@ Abstract.
We explain a Macaulay2 implementation of a construction, which appeared in [HolweckOeding23], of a graded algebra structure on the direct sum of a Lie algebra $\mathfrak{g}$ (typically $\mathfrak{sl}_{n}$) and a $\mathfrak{g}$-module (typically a subspace of an exterior algebra ${\textstyle\bigwedge^{\bullet}}\mathbb{C}^{n}$). We implement brackets, a Killing form, matrix representations of adjoint operators, and ranks of blocks of their powers.
@@ 1. Introduction
The Lie algebra $\mathfrak{g}=\mathfrak{sl}_{n}(\mathbb{F})$ is a matrix algebra over a field $\mathbb{F}$ with bracket $[A,B]=AB-BA$ for traceless matrices $A,B\in\mathbb{F}^{n\times n}$. The package ExteriorExtensions extends this algebra to a larger one $\mathfrak{a}=\mathfrak{g}\oplus M$ for a $\mathfrak{g}$-module $M$. At this time we have only implemented this for $M$ a sum of exterior powers of the underlying vector space $V=\mathbb{F}^{n}$. We say that the extension algebra has GJD (Jordan decomposition consistent with the $G$-action) if the $G=\text{Lie}(\mathfrak{g})$-action on $M$ is consistent with conjugation on adjoint operators $\operatorname{Ad}_{T}\colon\mathfrak{a}\to\mathfrak{a}$ (see [HolweckOeding23]) for $T\in\mathfrak{a}$. Our development is focused on making tools for studying elements $x\in M$. In particular, the conjugation invariants of an adjoint operator $\operatorname{Ad}_{x}$, such as the rank, the trace, and characteristic polynomial, are also $G$ invariants when the algebra $\mathfrak{a}$ has GJD. In addition, the adjoint operators inherit a block structure from the grading on $\mathfrak{a}$. This package enables one to compute these invariants for many situations of interest in particular for the invariant theory of tensors, and may be useful for applications in quantum information and algebraic statistics, as well as for the study of geometric invariants of special varieties like hyper Kähler manifolds, K3 surfaces and more. We explained these connections in [HolweckOeding23].
The main functions of the package are the following: buildAlgebra, which constructs a new type of graded ring, called an ExteriorExtension, that includes products implemented as bracket, and adjoint operators ad. These rings represent the algebra $\mathfrak{a}$. The package computes standard matrix representations of adjoint operators on $\mathfrak{a}$, and facilitates the tabulation of the ranks of blocks of powers of adjoint operators via blockRanks. We also compute the KillingMatrix, which may help identify the algebra $\mathfrak{a}$.
The remainder of the introduction provides a brief mathematical account of each of these objects. In Section 2 we provide examples of the functionality of the package, specifically these functions, and comment on the meaning of the computations. We note that M2 starts numbering by 0 by default, whereas we often start numbering at 1, and we hope that this shift between explanations and implementations is not too confusing for the reader.
@@ 1.1. Exterior algebras
The exterior algebra on a vector space $V$ is the direct sum of exterior powers ${\textstyle\bigwedge^{\bullet}}V=\bigoplus_{k\geq 0}{\textstyle\bigwedge^{k}}V$, with product given by concatenation, which respects a $\mathbb{Z}$-grading. The exterior algebra on $V$ is Artinian when $V$ is finite-dimensional. However, we can obtain a $\mathbb{Z}_{n}$-grading if $n=\dim V$ by choosing a volume form $0\neq\Omega\in{\textstyle\bigwedge^{n}}V$ and utilizing the correspondingly defined Hodge star $\star$ operation to replace products that would have degree greater than $n$. We explain this more fully in Section 1.2.1.
The vector space $V$ carries the standard group action by $G=\operatorname{SL}(V)$ and Lie algebra action by $\mathfrak{g}=\mathfrak{sl}(V)$. These actions are induced up to the exterior powers, making any direct sum of exterior powers $G$ (or $\mathfrak{g}$)-modules.
@@ 1.2. Extending the Lie algebra by appending an exterior algebra
Now fix positive integers $k$ and $n=\dim V$. The exterior algebra ${\textstyle\bigwedge^{\bullet}}V=\bigoplus_{k\geq 0}{\textstyle\bigwedge^{k}}V$ is graded by degree, however one obtains a (possibly) different grading by declaring ${\textstyle\bigwedge^{j}}V$ to be the grade-1 piece, and defining grading on the rest of the graded pieces inductively. More specifically, consider the graded vector space
$\mathfrak{a}=\mathfrak{g}\oplus{\textstyle\bigwedge^{k}}V\oplus{\textstyle\bigwedge^{2k}}V\oplus\dots=\mathfrak{g}\oplus\bigoplus_{ik\not\equiv 0\mod n}({\textstyle\bigwedge^{ik}}V),$
where we replace the scalars from the exterior algebra with $\mathfrak{a}_{0}=\mathfrak{g}$ and $\mathfrak{a}_{i}={\textstyle\bigwedge^{ik\mod n}}V$ for each $i$ from $i=1$ up to but not including the first $i$ such that $ik\equiv 0\mod n$. Note that if $k$ divides $n$ we are appending part of a sub-algebra of the exterior algebra of the parts whose degrees are multiples of $k$. For example, when $k=3$, and $n=9$ we construct
$\mathfrak{a}=\mathfrak{sl}_{9}\oplus{\textstyle\bigwedge^{3}}\mathbb{F}^{9}\oplus{\textstyle\bigwedge^{6}}\mathbb{F}^{9},$
but when $k=3,n=10$ we construct
$\mathfrak{a}=\mathfrak{sl}_{10}\oplus{\textstyle\bigwedge^{3}}\mathbb{F}^{10}\oplus{\textstyle\bigwedge^{6}}\mathbb{F}^{10}\oplus{\textstyle\bigwedge^{9}}\mathbb{F}^{10}\oplus{\textstyle\bigwedge^{2}}\mathbb{F}^{10}\oplus{\textstyle\bigwedge^{5}}\mathbb{F}^{10}\oplus{\textstyle\bigwedge^{8}}\mathbb{F}^{10}\oplus{\textstyle\bigwedge^{1}}\mathbb{F}^{10}\oplus{\textstyle\bigwedge^{4}}\mathbb{F}^{10}\oplus{\textstyle\bigwedge^{7}}\mathbb{F}^{10}.$
This vector space $\mathfrak{a}$ is constructed by the function buildAlgebra and the pieces that comprise it are accessed by LieAlgebra (for $\mathfrak{a}_{0}$) and appendage (for $\mathfrak{a}_{>0}$). More specifically, we construct $\mathfrak{a}_{0}=\mathfrak{sl}(V)$ by making a common choice basis for the traceless $n\times n$ matrices, and we access $\mathfrak{a}_{>0}$ by calling a ring with skew-commuting variables, $e_{i}$. One can access the bases we choose for these algebras by bases, which produces a hash table with key i calling the basis of $\mathfrak{a}_{i}$.
Already the graded vector space space $\mathfrak{a}$ has brakets (products) defined for $\mathfrak{a}_{i}\times\mathfrak{a}_{j}\to\mathfrak{a}_{i+j}$ when $1<i,j,i+j<n$ (the skew-commuting product on the ring $\mathbb{F}[\mathbb{F}^{n}]$). In the next section, we explain how to extend the definition of the bracket to the entire graded vector space in an equivariant fashion.
One of the main results from [HolweckOeding23] tells precisely when it is possible to make $\mathfrak{a}$ into a graded algebra respecting the group structure.
Theorem 1.1.
[HolweckOeding23, Theorem 3.18] The vector space $\mathfrak{a}=\mathfrak{sl}_{n}\oplus\bigoplus_{k=1\ldots n-1}{\textstyle\bigwedge^{k}}\mathbb{C}^{n}$ has a $\mathbb{Z}_{n}$-graded algebra structure with a Jordan decomposition consistent with the $G=\operatorname{SL}(V)$-action. There is a unique (up to scale) equivariant bracket product that agrees with the $\mathfrak{g}$-action on each $\mathfrak{g}_{i}$. If $n=2k$, the equivariant bracket must satisfy the property that the restriction to ${\textstyle\bigwedge^{k}}\mathbb{C}^{n}\times{\textstyle\bigwedge^{k}}\mathbb{C}^{n}\to\mathfrak{g}$ must be commuting when $k$ is odd and skew-commuting when $k$ is even. For any $k$ such that $2k\neq n$ the bracket ${\textstyle\bigwedge^{k}}\mathbb{C}^{n}\times{\textstyle\bigwedge^{k}}\mathbb{C}^{n}\to{\textstyle\bigwedge^{2k\mod n}}\mathbb{C}^{n}$ must be skew-commuting when $k$ is odd and commuting when $k$ is even.
This result here is that a natural equivariant bracket exists, so it makes sense to try to implement it in the first place, and moreover, since the structure tensor $B$ for the algebra $\mathfrak{a}$ is essentially unique, there aren’t many choices in its formulation. In particular, the only choices we have are choices of scalar multiples of the blocks $B_{i,j,i+j}$ of the tensor. We implement that product as a method function in M2 called bracket. Next, we give a detailed description of the essentially unique structure tensor for this algebra.
1.2.1. The structure tensor defining the bracket
If $V$ is a vector space over a field $\mathbb{F}$, we let $V^{*}$ denote the dual vector space of $\mathbb{F}$-linear functionals. Following the perspective of [LandsbergTensorBook] a tensor $T\in V_{1}\otimes\cdots\otimes V_{m}$ can be equivalently viewed in many ways, such as a linear functional $T\colon V_{1}^{*}\otimes\cdots\otimes V_{m}^{*}\to\mathbb{F}$, as a linear map $T\colon V_{1}^{*}\otimes\cdots\otimes\widehat{V^{*}_{i}}\otimes\cdots\otimes V_{m}^{*}\to V_{i}$ (where $\widehat{\cdot}$ denotes omission), or as a multilinear map $T\colon V_{1}^{*}\times\cdots\times V_{m}^{*}\to\mathbb{F}$.
@@ 1.2.1. The structure tensor defining the bracket
Recall that a structure tensor $B\in\mathfrak{a}^{*}\otimes\mathfrak{a}^{*}\otimes\mathfrak{a}$ representing the bilinear product $[\;,\;]\colon\mathfrak{a}\times\mathfrak{a}\to\mathfrak{a}$ is equivalent information to the algebra structure on $\mathfrak{a}$. We denote by $B_{i,j,k}\in\mathfrak{a}_{i}^{*}\otimes\mathfrak{a}_{j}^{*}\otimes\mathfrak{a}_{k}$ the graded pieces of $B$. To respect the grading we must require that $B_{i,j,k}=0$ whenever $i+j\not\equiv k\mod n$.
@@
For $B_{0,0,0}$ we have the standard skew-commuting bracket on $\mathfrak{sl}_{n}$, namely $[A,B]=AB-BA$. For $B_{0,j,j}$ (and for $B_{j,0,j}$) with $j>0$ we use the standard Lie algebra action of $\mathfrak{sl}_{n}$ on the fundamental representation ${\textstyle\bigwedge^{jk}}\mathbb{C}^{n}$, namely for $A\in\mathfrak{sl}_{n}$ and $T\in{\textstyle\bigwedge^{jk}}V$ we take $[A,T]:=A.T$, where . denotes the Lie algebra action. Similarly $[T,A]=-A.T$.
For $i>0$ we have two possibilities for $B_{i,j,(i+j)\mod n}$ depending on whether or not $(i+j)\mod n=0$. When $(i+j)\equiv 0\mod n$ we separate two further cases depending on whether $i\leq n/2$. When $0<i+j<n$ we use the standard product in the exterior algebra as the bracket.
If $i+j\equiv 0\mod n$, for $S$ of degree $i\leq n/2$ and $T$ of degree $j>n/2$ we utilize the Hodge star $\star T$ to change degrees and compute as follows. We declare an orthonormal basis $e_{1},\ldots,e_{n}$ of $V=\mathbb{F}^{n}$ and note that the Hodge star is determined by a volume form which we choose to be $\Omega=e_{1}\wedge\cdots\wedge e_{n}$. The standard Hodge star operator for $T\in{\textstyle\bigwedge^{j}}V$ producing $\star T\in{\textstyle\bigwedge^{n-j}}V$ is implemented as follows:
$\star T=\sum_{\alpha\in\binom{[n]}{j}}\text{sgn}(\alpha)\left(\big{(}\prod_{i\in\alpha}e_{i}\big{)}\lrcorner T\right)\prod_{i\in\alpha^{c}}e_{i},$
@@
We have chosen a global scalar to calibrate and match the outcomes of [Vinberg-Elashvili] (see the end of example 4) which considered the case $\mathfrak{sl}_{9}\oplus{\textstyle\bigwedge^{3}}\mathbb{C}^{9}\oplus{\textstyle\bigwedge^{3}}\mathbb{C}^{9}$ as a model for the exceptional Lie algebra $\mathfrak{e}_{8}$. In summary, the map that takes two elements of the exterior extension to a matrix in $\mathfrak{sl}_{n}$ is the following:
$\begin{matrix}\mathfrak{a}_{i}\times\mathfrak{a}_{j}&\to&\mathfrak{a}_{0}\\ (S,T)&\mapsto&[S,T]=-k\pi\left(\left(\frac{\partial S}{\partial e_{p}}\lrcorner\frac{\partial\star T}{\partial e_{q}}\right)_{1\leq p,q\leq n}\right)\end{matrix}\quad\quad\text{when }i+j=n\text{ and }j>n/2.$
For $i\geq n/2$ and $j<n/2$ we reverse the role of the Hodge dual and proceed as in the previous paragraph.
Finally, when $i+j>n$ we use Hodge star to convert to a product that we have already computed. If $S$ has degree $i>n/2$ or $T$ has degree $j>n/2$ then we take $[S,T]=\star[\star S,\star T]$ and use the bracket previously defined for $i+j\leq n$.
Finally, we remark that the products that have degree sum $i+j>n$ on the level of basic elements $S=e_{I}$ and $T=e_{J}$ essentially result in $[S,T]\sim e_{I\cap J}$, however one has to pay attention to signs. Likewise, when the degree sum is $i+j<n$ on the level of basic elements $S=e_{I}$ and $T=e_{J}$ essentially results in $[S,T]\sim e_{I\cup J}$, however again one has to pay attention to signs.
@@ 1.2.2. Adjoint Operators
The standard matrix representation of the adjoint operator on $\mathfrak{a}$ defined by $\operatorname{Ad}_{x}(y)=[x,y]$ is constructed by choosing an ordered basis of $\mathfrak{a}$, and expressing the results of $\operatorname{Ad}_{x}(y)$ for each basis vector $y$ in terms of the standard basis as the corresponding column. One caveat is that we have to choose a basis of $\mathfrak{sl}_{n}$, and there is not a "standard" choice for the basis of the Cartan subalgebra of traceless diagonal matrices. We choose a common one $\{E_{ii}-E_{i+1,i+1}\mid i=1\ldots n-1\}$. Then we have to solve a system of equations to find the coordinates for the cartan rather than just extracting the entries of the matrix. Solving this system of equations can depend on the field $\mathbb{F}$ because we invert a matrix over $\mathbb{F}$ to find the solution.
Not that the adjoint operator is a linear map $\operatorname{Ad}_{x}\colon\mathfrak{a}\to\mathfrak{a}$, which is represented by a (large) matrix. The structure tensor is graded so that $B_{i,j,k}\colon\mathfrak{a}_{i}\times\mathfrak{a}_{j}\to\mathfrak{a}_{k}$. The map respects the grading on $\mathfrak{a}$ implies that $B_{i,j,k}=0$ unless $k\equiv i+j\mod n$. This, in turn, induces a block structure on the matrix representing $\operatorname{Ad}_{x}$ since $\operatorname{Ad}_{x}(y)=\bigoplus_{i,j,k}(B_{i,j,k}(x,y))$. For example, in the 2-graded case, with $x\in\mathfrak{a}_{1}$ we have the blocking
$\operatorname{Ad}_{x}=\left(\begin{smallmatrix}0&*\\ *&0\end{smallmatrix}\right),$
where the 00’s and *’s respectively represent block matrices, and column blocks (and row blocks) are respectively of sizes $(\dim\mathfrak{a}_{0}$, $\dim\mathfrak{a}_{1})$. In the 3-graded case, with $x\in\mathfrak{a}_{1}$ we have the blocking
$\operatorname{Ad}_{x}=\left(\begin{smallmatrix}0&0&*\\ *&0&0\\ 0&*&0\end{smallmatrix}\right),$
with column blocks (and row blocks) respectively of sizes $(\dim\mathfrak{a}_{0}$, $\dim\mathfrak{a}_{1},\dim\mathfrak{a}_{2})$.
The ranks of the blocks of $\operatorname{Ad}_{x}$ are invariants of $x$, as are the ranks of blocks of the powers of $\operatorname{Ad}_{x}$. We collect these ranks in a table and display them with a function prettyBlockRanks.
We note that we implement the ranks of powers of blocks “under the hood” by extracting blocks of $(\operatorname{Ad}_{x})^{k}$ as submatrices of the appropriate sizes (determined by the dimensions of the graded pieces, and the respective ordered bases we constructed to carry out our computations) and computing their ranks, and moreover, we compute $(\operatorname{Ad}_{x})^{k}$ sequentially. We note that it is, in theory, possible to make a generic adjoint operator since the map $\mathfrak{a}\to\operatorname{End}\mathfrak{a}$ is linear, and that would produce $\operatorname{Ad}_{x}$ as a matrix of linear forms on $\mathfrak{a}$. Computing with a large matrix of linear forms can quickly become costly computationally, so we prefer to only work with specific numerical instances, where computing powers and ranks of powers is computationally inexpensive.
The Jacobi identity is stated via adjoint operators as $[\operatorname{Ad}_{x},\operatorname{Ad}_{y}]=\operatorname{Ad}_{[x,y]}$ [FultonHarris]. We can check whether this identity holds for the algebras we construct either on random elements of each grade, or on bases.
@@ 1.2.3. The Killing form
1.2.4. On fields of definition
This package was written with the fields $\mathbb{C}$ or $\mathbb{R}$ in mind because of our knowledge of results in representation theory over the real and complex numbers. However, much of the functionality of M2 is better suited for exact arithmetic, over $\mathbb{Q}$ or $\mathbb{Z}_{p}$, for instance. In one place the package attempts to find an expression of a matrix in terms of a specially chosen basis of $\mathfrak{sl}_{n}$, which we carry out by solving a system of equations. Solving this linear system inverts a matrix (the locally defined matrix Qmat in the package), and in M2 this operation has different behavior based on the ground field. At this time of development, the author has not tested behavior over other rings or fields, and so the functions are implemented over the M2 type Ring, and M2 will throw an error when imprecise fields like $\mathbb{R}$ or $\mathbb{C}$ are passed to the buildAlgebra command. It is still possible to have some of this functionality, and to do this is typically best to start working over an exact field, and only lift to $\mathbb{R}$ or $\mathbb{C}$ when necessary or when it makes sense to do so. By not working over inexact fields until necessary we reduce the chances for numerical errors.